OpenStudy (anonymous):

What is the slope-intercept form of the function described by this table? (Im trying to make a rough graph with keyboard keys) x|1| 2| 3| 4| ----------- y|1|-2|-5|-8| y = Blank X + Blank

OpenStudy (anonymous):

so first solve for the slope. take to of the sets of coodinates and solve them like this (y2 -y1)/(x2-x1)

OpenStudy (anonymous):

and that will solve for m. to solve for b, which is the y intercept, take a point and graph it with the slope.

OpenStudy (campbell_st):

what do the consecutive y values change by..? that will be the slope

OpenStudy (anonymous):

campbell*

OpenStudy (campbell_st):

then there are 2 choices for the 2nd blank... when you know m..... substitute a point from the table, say (1, 1) then you can find the 2nd blank.

OpenStudy (campbell_st):

so what is y changing by...?

OpenStudy (anonymous):

im back i was doing some stuff but thanks so i got the Slope m = -3. Now how would i find the intercept?

OpenStudy (anonymous):

@Dfoeltz btw not sure if thats how you tag.

OpenStudy (anonymous):

@Dfoeltz y = -3X + B then i did -8 = -3 * 4 + b then i got -8 = -12 + b then do i add 12 to both sides?

OpenStudy (campbell_st):

great... so the equation is y = -3x + b if you use the table add an extra column x: 0 : 1 : 2: 3 : y: : 1 : -2: -5: use y = 1 and subtract the slope 1 - (-3) where do you start the pattern... the alternate method is to use a point and substitute it into y = -3x + b choose 2, -3 so -3 = -3*2 + b solve for b.... you can choose any point and get the same answer

OpenStudy (anonymous):

Thanks for all of your guys help.